Machine cost = 5 lack
Salvage value = 54994.35
Find book value of machine after 5 years and also find accumulated depreciation after nine years .
By using db method
1
Expert's answer
2020-09-30T13:17:03-0400
"Solution"

Straight-line Depreciation rate"=\frac15=0.2=20\%"

Declining Balance Rate"=2\times20\%=40\%"

Depreciation"=40\%\times500000=200,000"

Then, Calculate the depreciation of the asset for the second year of it's life.

Declining rate is still "40\%"

"Book\ value=Cost-Accumulated\ Depreciation\\
=500,000-200,000\\
=300,000\\
Depreciation=40\%\times300,000\\
=120,000"

Third Year.

"Book\ value=500,000-200,000-120,000\\
=180,000\\
Depreciation=40\%\times180,000\\
=72,000"

Fourth Year.

"Book\ value=500,000-200,000-120,000-72,000\\
=108,000\\
Depreciation=40\%\times108,000\\
=43,200"

Fifth Year.

"Book\ value=500,000-200,000-120,000-72,000-43,200\\
=64,800\\
Depreciation=40\%\times64,800\\
=25,920"

The depreciation calculated above will decrease the Book value of the asset below it's estimated residual value("64,800-25920=38880<54994.35" ).Therefore, depreciation would only be allowed up to the point where the Book value=Salvage value. This depreciation expense of"9,805=(64,800-54,994.35) is\ allowed." Same procedure will be replicated for 9years.